An unknown metal is found to have a density of 7.8748g/cm3and to crystallize in a body-centered cubic lattice. The edge of the unit cell is found to be 0.31627nm. Calculate the atomic mass of the metal.
Answer #1

Density \rho = 7.8748 g/cm3

For body centre Z = 2

\rho = Z x M / a3 x N0

7.8748 = 2 x M / [(0.31627 X 10-7)3 x 6.02 x 1023]

M = (7.8748 x( 0.31627 X 10-7)3 x 6.02 x 1023)/2 = 0.7498 x 102 g/mol = 74.98 g/mol
Answer #2

each cell=(0.31627*10^-9) - cell V. atomic mass= 7.8748g/cm3* V Cm3/2*6.023*10^23=

7.49957E+14 atoms/mole